1. Find the radius of a circle with circumference of 20pi.
2.Find the radius of a circle with the circumference of 64pi
3.If the area of a circle is 9pi, find the radius.
4.If the area of a circle is 64 pi, find the circumference.


help much appreciated.

Found 2 solutions by stanbon, checkley71:
Answer by stanbon(48558)

1. Find the radius of a circle with circumference of 20pi.
2.Find the radius of a circle with the circumference of 64pi
3.If the area of a circle is 9pi, find the radius.
4.If the area of a circle is 64 pi, find the circumference.
#1. Circumference = 2(pi)r=20(pi)
r=10
#2. 2(pi)r=64(pi)
f = 32
#3. Area = (pi)r^2 = 9(pi)
r^2=9
r = 3
#4 (pi)r^2=64(pi)
r^2=64
r = 8
Cheers,
Stan H.


Answer by checkley71(8403)

CIR=2*PI*R OR 20*PI=2*PI*R DIVIDING BOTH SIDES BY 2*PI WE GET R=10
CIR=2*PI*R OR 64*PI=2*PI*R DIVIDING BOTH SIDES BY 2*PI WE GET R=32
AREA=PI*R~2 OR 9*P1=PI*R~2 DIVIDING BOTH SIDES BY PI WE GET R~2=9 OR R=3
AREA=PI*R~2 OR 64*PI=PI*R~2 DIVIDING BOTH SIDES BY PI WE GET R~2=54 OR R=8